OpenStudy (anonymous):

a rectangular table is six times as long as it is wide. if the area is 150ft^2 find the length and width of the table

jhonyy9 (jhonyy9):

A=l*w l=6w A=6w*w 150=6w*w 150/6 =w*w 25=w*w 5*5=w*w w=5 l=6*5=30

jhonyy9 (jhonyy9):

right ?

jhonyy9 (jhonyy9):

so if yes please push the good answer button

OpenStudy (anonymous):

thanks